# Causal Network Recovery in Perturb-seq Experiments Using Proxy and Instrumental Variables

### v1

@@ 4.1 Estimation of the descendant sets
We estimate the variance of $\hat{\tau}^{*}_{jk}$ using a robust (sandwich) estimator. We refer to Section LABEL:supp-sec:sandwich for details on the variance estimator $\hat{V}_{\tau^{*}_{jk}\tau^{*}_{jk}}$ for different choices of the expression models. With the robust variance estimator, the Wald statistic for testing $H_{0jk}\!:\tau_{jk}^{*}=0$ can be constructed as
@@
which yields asymptotically valid inference even if the other parameters are misspecified, as long as the mean function (5) is correctly specified (white1982maximum). i.e., we have
$z_{\tau^{*}_{jk}}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1),$ (8)
as $N\rightarrow\infty$, under $H_{0jk}$.
We collect the p-values
@@
across $j,k\in[p]$ for $j\neq k$ and employ p-value adjustment based on BH, (benjamini1995controlling) FDR control at certain pre-specified $\alpha\in(0,1)$.
If we choose to rely on classical influentiality assumption (tian2013causal, yang2018characterizing), we can finalize the descendant search by defining $\hat{des}(j)\coloneqq\{m\in[p]/\{j\}|p_{jk}<\alpha_{adj}\}.$
@@ 4.2 Estimation of the parent sets
Note that the proposed two-stage regression framework is closely related in spirit to the classical two-stage least squares (TSLS) approach with valid instrumental variables in linear models. Here, ${\mathbf{D}}_{\mathrm{anc}(j)}$ participates in the entire first-stage regressions across $k$ and serves as a valid instrumental variable for $j$. Specifically, it satisfies: (i) conditional independence from unobserved confounders, $U\perp\!\!\!\!\perp{\mathbf{D}}_{anc(j)}\mid X$; (ii) the exclusion restriction, as there is no direct causal path from ${\mathbf{D}}_{anc(j)}$ to $\mu_{j}$; and (iii) relevance, since $D_{k}$ directly perturbs $\mu_{k}$ individually for $k\in anc(j)$.
Here, the uncertainty of $\hat{\eta}_{k}({\mathbf{D}},X)$ needs to be taken into account for a correct quantification of the standard error of $\hat{\theta}_{jk}$ estimated from QMLE. We consider the regression coefficients $\hat{\theta}_{jk}$ for each $\hat{\eta}_{k}$ based on a GLM, e..g, Poisson GLM, and quantify the standard error of $\hat{\eta}_{k}$ based on the murphy2002estimation standard error adjustment for estimated regressors.
Specifically, following murphy2002estimation, we correct the standard error of $\hat{\theta}_{jk}$ to account for the sampling uncertainty in the estimated regressors $\hat{\eta}_{k}({\mathbf{D}},X)$, each obtained from a separate first-stage GLM fit (11). For each ancestor $k\in anc(j)$, let $\hat{V}_{\xi_{k}\xi_{k}}$ be the sandwich covariance of $\hat{\xi}_{k}$ from this fit, and define $\hat{V}_{\xi\xi}=\mbox{block-diag}(\hat{V}_{\xi_{k}\xi_{k}}:k\in anc(j))$ as the block-diagonal covariance across all ancestors.
Note that we estimate the parent coefficients $\theta_{jk}$ via the second-stage GLM in (12). Let $s_{i}(\theta_{j},\xi)$ denote its score function, where $\theta_{j}$ collects all second-stage regression coefficients for $Y_{ij}$ with regressors $\{\hat{\eta}_{k}({\mathbf{D}}_{i},X_{i})\}_{k\in anc(j)}$, and $\xi$ denotes the concatenation of $\xi_{k}$ over $k\in anc(j)$. Then the Murphy–Topel asymptotic covariance of $\hat{\theta}_{j}$ is
@@
where $\hat{A}_{\theta\theta}=-\tfrac{1}{N}\sum_{i}\partial s_{i}/\partial\theta_{j}^{\top}$, $\hat{B}_{\theta\theta}=\tfrac{1}{N}\sum_{i}s_{i}s_{i}^{\top}$, and $\hat{A}_{\theta\xi}=[\hat{C}_{k}]_{k\in anc(j)}$ collects the cross-derivative blocks $\hat{C}_{k}=-\tfrac{1}{N}\sum_{i}\partial s_{i}/\partial\xi_{k}^{\top}$. The additional term $\hat{A}_{\theta\xi}\hat{V}_{\xi\xi}\hat{A}_{\theta\xi}^{\top}$ captures the variability propagated from all first-stage regressions used to estimate $\hat{\eta}_{k}({\mathbf{D}},X)$, ensuring asymptotically valid inference for each $\hat{\theta}_{jk}$ even when all regressors $\hat{\eta}_{k}$, $k\in anc(j)$, are estimated rather than observed. Then, under standard regularity conditions in murphy2002estimation for those QMLEs, we have
@@
as $N\rightarrow\infty$, providing asymptotically valid Wald tests for $\theta_{jk}$ (Fig. LABEL:supp-fig:mt_se). The corresponding p-values are used in online FDR control (zrnic2020power) in our final DAG search algorithm. The Gaussian approximation for the edge coefficients is suitable for Perturb-seq data, where the number of genes $p$ generally satisfies $p\ll N$. Although we do not employ this strategy in our framework, one may also consider fitting a regularized GLM to address potential high-dimensional settings.
@@
When the parent information is perfect for all $j\in V(G)$, the full DAG is directly obtainable by concatenating the parent nodes of each node together. However, in finite-sample settings, estimation errors in parent identification may occur, and naive concatenation of the estimated parent sets can inadvertently introduce cycles into the resulting graph. Here, we introduce an algorithm (Algorithm 1) that is guaranteed to yield a DAG even with only finite samples available.
@@
Then, the algorithm builds a DAG starting from the sink nodes with no descendants and sequentially progresses toward the root nodes with the largest number of descendants, connecting each previous node to the next with an edge. In this way, it is guaranteed to output a directed, acyclic graph, namely a DAG.
The online FDR control is needed in Step 3 of the algorithm since the number of the tests deppends on the estimated ancestry set $\hat{anc}(j)$. Based on the p-values $p_{\theta_{jk}}=1-\Phi(z_{\theta_{jk}})$ across $k\in\hat{anc}(j)$, collected for each response gene $j$, we employ the online batch FDR control by zrnic2020power for a sequential adjustment of sets of the p-values across $j$. We particularly need this online adjustment, as the set of p-values that arrives next depends on what is selected previously based on an DAG search algorithm that we present in the next section. We define the estimated parent set as
@@
where $\alpha_{j}$ is an online adjusted p-value cut-off updated for each $j$ by zrnic2020power.
@@ 5 Simulation
For a simple illustration, we generated a Perturb-seq dataset of $N=8000$ and $p=8$ that involves unmeasured confounder $U$ and observed confounder $X$ in the data generative model (Fig. 2a). (Section LABEL:supp-sec:simul_DAG for details and Fig. LABEL:supp-fig:Illust_example_medp for a larger $p$ scenario). Each perturbation is represented by approximately 350 cells, while about 5,000 cells serve as controls. For the experiment, we drop $Y_{7}$ and $Y_{8}$ from the analyses and apply ARGEN to the simulated datasets replicated 100 times. For baseline comparisons, we considered INSPRE (brown2025large) and GLM with naive proxy approximation of the true expression.
@@ 6 Identification of Causal Essential Gene Networks from K562 Cells with Perturb-seq Experiments
We present results of detailed analysis using the proposed method ARGEN to learn the causal gene networks among sets of essential genes based on the Perturb-seq data, which treated the K562 cells with CRISPR interference (CRISPRi) intervention to reduce the expression levels of essential genes (replogle2022mapping). These genes involve in transcription and chromatin Regulation RNA processing, splicing and RBP regulation, proteasome and ubiquitin–proteostasis, cell cycle, DNA replication and genome stability, translation initiation and elongation factors, ribosomal proteins, mitochondrial translation, respiratory chain and metabolism, cytoskeleton, motor and structural components, and vesicle trafficking and membrane transport etc. Because each of the essential genes is perturbed directly by CRISPRi in K562 cells, the resulting perturb-seq data provide an interventional framework for estimating causal networks among these genes, representing a context-specific causal effect in this cellular system. While such effects may reflect both direct and mediated regulatory relationships, the proposed approach is expected to yield a biologically interpretable causal gene network among core essential regulators in K562.
@@ 6.1 Evaluation of ARGEN intra-chromosomal causal networks from 3D genome and epigenomic datasets
We first start with analyzing the Perturb-seq data in a chromosome-specific manner and provide evaluations on the causal network learned from the data. We focused on 874 perturbations that had enough number of cells and CRISPRi signal on the targeted genes. For further details on the selection of such genes, refer to Section LABEL:supp-sec:processing_implementation. For each of the 23 human chromosomes, we collected the genes that reside within the same chromosome and applied ARGEN to infer a gene influence network. This resulted in sets of DAGs with around 40 genes and 100 edges per each chromosome on average.
We first implemented an initial sanity check for the directed edges in the DAGs. We checked whether the directional relationships between pairs of genes are indeed causal, instead of being associational. Specifically, as the data is from CRISPRi experiment, where the guide RNA is designed to reduce the expression level of the targeted gene, we would expect to have $\tau_{k}<0$ in the SEM (2). Thus, if the estimated edge coefficient $\hat{\theta}_{jk}<0$ for a parent gene $k$ and a child gene $j$, we would expect the expression level of gene $j$ to increase under an intervention on gene $k$; conversely, if $\hat{\theta}_{jk}>0$, we would expect the expression level of gene $j$ to decrease. We first checked whether such a relationship holds for each target gene. As an illustrative example, we collected the entire parents of the gene MCM3 in chromosome 6 (Fig. 3a) and checked the relationship. It can be observed from Fig. 3b that the perturbation on each of the parent genes of MCM3 result in significant increase of MCM3 expression whenever the coefficient $\hat{\theta}_{jk}<0$ and significant decrease of MCM3 expression whenever the $\hat{\theta}_{jk}>0$, aligning with our expectation for CRIPSRi experiments. On the other hand, when MCM3 was perturbed reversely, the expression changes of its parent genes neither exhibit this inverse relationship nor reach comparable levels of significance.
@@
We then checked whether such a result holds for all the edges in the 23 chromosome specific DAGs and 98% of the ARGEN coefficients $\hat{\theta}_{jk}$ show negative association with the intervention treatment effect, learned from perturbing a parent of a gene and evaluating the expression change of the gene (Fig. 3e). In comparison, only around 55% of the ARGEN coefficients showed such relationship with the reversed treatment effect, where a child gene is intervened and the change in expression of a parent node of the child was measured, indicating the sign match of nearly random chance (50%). These results reveal that the directionality of the parents to child edge is likely to be causal instead of associational.
@@
Next, we checked whether ARGEN DAGs show biological support from 3D genomics and epigenomics. We first investigated the genomic distance pattern of the DAG edges. We specifically gathered the entire $\log_{10}(\mbox{p-values})$ of $\hat{\theta}_{jk}$ and investigated how the edge signal associates with the genomic distance between parent and child genes. We observe that the regulatory elements far apart from the target genes are generally less supported by ARGEN p-values (Fig. 4a), aligning with previous evidences on the genomic distance of gene regulatory elements from the target genes (fulco2019activity).
@@
On top of genomic distance-based evaluation, we further sought for supports from physical proximity information captured from the same system. We specifically explored the actual physical distance structure of the ARGEN edges based on an external Hi-C data of K562 cells (rao20143d). We first observed that the gene pairs with no Hi-C contact had significantly ($p<2\times 10^{-5}$) lower edge signals than the ones with non-zero Hi-C contacts. In addition, the gene pairs with stronger regulatory signals by ARGEN tend to be supported with higher normalized Hi-C contacts than the ones with less signals ($p<3\times 10^{-6}$, Fig. LABEL:supp-fig:hic_contact).
@@
A/B compartments, which form large genomic territories with relatively high (A) or low (B) gene expression compared to other territories, constitute a class of chromatin domain compartmentalization (lieberman2009comprehensive). Consistent with recent findings (zufferey2021systematic) that co-regulation within chromatin domains is more prevalent among lowly expressed or less efficiently transcribed genes, as well as evidence that closed chromatin regions exhibit elevated co-methylation across loci (fortin2015reconstructing), we observed that the overall edge signal from ARGEN, with compartments learned from Hi-C data, was significantly higher in the B compartment than in the A compartment ($p<5\times 10^{-3}$). Furthermore, among the gene pairs with Transcription Start Site (TSS) distance less than 2 Mega base pairs (Mb), we checked how many of the parent genes are within the same Topologically Associating Domain (TAD), a contiguous regions of genome within which chromatin interactions are enriched (dixon2012topological). We investigated the proportion of such regulatory genes in the gene pairs called as significant by ARGEN’s online FDR procedure in comparison to those in the gene pairs called as not significant. This comparison yielded that those genes within ARGEN significant edges show significantly higher proportion ($p<10^{-3}$) than the other set of genes (Fig. 4c). Furthermore, those regulatory genes called by ARGEN showed significantly higher proportion ($p<10^{-3}$) than the ones called by INSPRE and Naive GLM, while the proportion of the baseline methods showed no significant difference compared to that of the gene pairs randomly selected (Fig. 4c).
@@
Finally, we checked whether the gene pairs called from ARGEN showed enriched ChIP-seq signal of binding of Transcription Factors relevant to K562 and one of the Histone marks (H3K27ac) studied in gasperini2019genome, barry2024robust. We specifically quantified the proportion of edges involving enriched ChIP-seq signal in both of the gene promoter regions, defined as 1Kb upstream region from the TSS, and compared the proportion against that of randomly selected gene pairs. We observed that six (yellow colored panels in Fig. 4d) out of eight ChIP-seq targets showed enrichment in the ARGEN edges. In comparison, half of the ChIP-seq targets showed lack of enrichment in the INSPRE edges.
Collectively, these evaluation results provide biological support of ARGEN DAG edges in terms of causal relationship among the genes.

### v2

@@ 4.1 Estimation of the descendant sets
We estimate the variance of $\hat{\tau}^{*}_{jk}$ using a sandwich estimator. We refer to Section LABEL:supp-sec:sandwich for details on the variance estimator $\hat{V}_{\tau^{*}_{jk}\tau^{*}_{jk}}$ for different choices of the expression models. With the variance estimator, the Wald statistic for testing $H_{0jk}\!:\tau_{jk}^{*}=0$ can be constructed as
@@
which yields asymptotically valid inference even if the other parameters are misspecified, as long as the mean function (5) is correctly specified (White 1982), i.e., we have $z_{\tau^{*}_{jk}}\stackrel{{\scriptstyle d}}{{\rightarrow}}N(0,1),$ as $N\rightarrow\infty$, under $H_{0jk}$. We collect the p-values
@@
across $j,k\in[p]$ for $j\neq k$ and employ p-value adjustment based on BH(Benjamini and Hochberg 1995) FDR control at certain pre-specified $\alpha\in(0,1)$.
If we choose to rely on classical influentiality assumption (Tian and Pearl 2013, Yang et al. 2018), we can finalize the descendant search by defining $\hat{des}(j)\coloneqq\{m\in[p]/\{j\}|p_{jk}<\alpha_{adj}\}.$
@@ 4.2 Estimation of the parent sets
The proposed two-stage framework is closely related in spirit to the classical two-stage least squares approach with valid instrumental variables. Here, ${\mathbf{D}}_{\mathrm{anc}(j)}$ participates in the entire first-stage regressions across $k$ and serves as a valid instrumental variable for $j$. Specifically, it satisfies: (i) conditional independence from unobserved confounders, $U\perp\!\!\!\!\perp{\mathbf{D}}_{anc(j)}\mid X$; (ii) the exclusion restriction, as there is no direct causal path from ${\mathbf{D}}_{anc(j)}$ to $\mu_{j}$; and (iii) relevance, since $D_{k}$ directly perturbs $\mu_{k}$ individually for $k\in anc(j)$.
Here, the uncertainty of $\hat{\eta}_{k}({\mathbf{D}},X)$ needs to be taken into account for a correct quantification of the standard error of $\hat{\theta}_{jk}$ estimated from QMLE. We consider the regression coefficients $\hat{\theta}_{jk}$ for each $\hat{\eta}_{k}$ based on a GLM, e..g, Poisson GLM, and quantify the standard error of $\hat{\eta}_{k}$ based on the Murphy and Topel (2002) standard error adjustment for estimated regressors.
Specifically, following Murphy and Topel (2002), we correct the standard error of $\hat{\theta}_{jk}$ to account for the sampling uncertainty in the estimated regressors $\hat{\eta}_{k}({\mathbf{D}},X)$, each obtained from a separate first-stage GLM fit (10). For each ancestor $k\in anc(j)$, let $\hat{V}_{\xi_{k}\xi_{k}}$ be the sandwich covariance of $\hat{\xi}_{k}$ from this fit, and define $\hat{V}_{\xi\xi}=\mbox{block-diag}(\hat{V}_{\xi_{k}\xi_{k}}:k\in anc(j))$ as the block-diagonal covariance across all ancestors.
Here we estimate the parent coefficients $\theta_{jk}$ via the second-stage GLM in (11). Let $s_{i}(\theta_{j},\xi)$ denote its score function, where $\theta_{j}$ collects all second-stage regression coefficients for $Y_{ij}$ with regressors $\{\hat{\eta}_{k}({\mathbf{D}}_{i},X_{i})\}_{k\in anc(j)}$, and $\xi$ denotes the concatenation of $\xi_{k}$ over $k\in anc(j)$. Then the Murphy–Topel asymptotic covariance of $\hat{\theta}_{j}$ is
@@
where $\hat{A}_{\theta\theta}=-\tfrac{1}{N}\sum_{i}\partial s_{i}/\partial\theta_{j}^{\top}$, $\hat{B}_{\theta\theta}=\tfrac{1}{N}\sum_{i}s_{i}s_{i}^{\top}$, and $\hat{A}_{\theta\xi}=[\hat{C}_{k}]_{k\in anc(j)}$ collects the cross-derivative blocks $\hat{C}_{k}=-\tfrac{1}{N}\sum_{i}\partial s_{i}/\partial\xi_{k}^{\top}$. The additional term $\hat{A}_{\theta\xi}\hat{V}_{\xi\xi}\hat{A}_{\theta\xi}^{\top}$ captures the variability propagated from all first-stage regressions used to estimate $\hat{\eta}_{k}({\mathbf{D}},X)$, ensuring asymptotically valid inference for each $\hat{\theta}_{jk}$ even when all regressors $\hat{\eta}_{k}$, $k\in anc(j)$, are estimated rather than observed. Then, under standard regularity conditions in Murphy and Topel (2002) for those QMLEs, we have
@@
as $N\rightarrow\infty$, providing asymptotically valid Wald tests for $\theta_{jk}$ (Fig. LABEL:supp-fig:mt_se). The corresponding p-values are used in online FDR control (Zrnic et al. 2020) in our final DAG search algorithm. The Gaussian approximation for the edge coefficients is suitable for Perturb-seq data, where the number of genes $p$ generally satisfies $p\ll N$. Although we do not adopt this strategy, one could fit a regularized GLM to handle high-dimensional settings.
@@
When the parent information is perfect for all $j\in V(G)$, the full DAG is directly obtainable by concatenating the parent nodes of each node together. However, in finite-sample settings, estimation errors in parent identification may occur, and naive concatenation of the estimated parent sets can inadvertently introduce cycles into the resulting graph. Here, we introduce Algorithm 1 that is guaranteed to yield a DAG even with only finite samples available.
@@
Our proposed algorithm builds a DAG starting from the sink nodes with no descendants and sequentially progresses toward the root nodes with the largest number of descendants, connecting each previous node to the next with an edge. In this way, it is guaranteed to output a directed, acyclic graph, namely a DAG.
The online FDR control is needed in Step 3 of the algorithm since the number of the tests depends on the estimated ancestry set $\hat{anc}(j)$. Based on the p-values $p_{\theta_{jk}}=1-\Phi(z_{\theta_{jk}})$ across $k\in\hat{anc}(j)$, collected for each response gene $j$, we employ the online batch FDR control by Zrnic et al. (2020) for a sequential adjustment of sets of the p-values across $j$. We particularly need this online adjustment, as the set of p-values that arrives next depends on what is selected previously based on an DAG search algorithm that we present in the next section. We define the estimated parent set as
@@
where $\alpha_{j}$ is an online adjusted p-value cut-off updated for each $j$ per Zrnic et al. (2020).
@@ 5 Simulation
For a simple illustration, we generated a Perturb-seq dataset of $N=8000$ and $p=8$ that involves unmeasured confounder $U$ and observed confounder $X$ in the data generative model (Fig. 2a). (Section LABEL:supp-sec:simul_DAG for details and Fig. LABEL:supp-fig:Illust_example_medp for a larger $p$ scenario). Each perturbation is represented by approximately 350 cells, while about 5,000 cells serve as controls. For the experiment, we drop $Y_{7}$ and $Y_{8}$ from the analyses and apply ARGEN to the simulated datasets replicated 100 times. For baseline comparisons, we considered INSPRE (Brown et al. 2025) and GLM with naive proxy approximation of the true expression.
@@ 6 Identification of Causal Essential Gene Networks from K562 Cells with Perturb-seq Experiments
We present results of detailed analysis using the proposed method ARGEN to learn the causal gene networks among sets of essential genes based on the Perturb-seq data, which treated the K562 cells with CRISPR interference (CRISPRi) intervention to reduce the expression levels of essential genes (Replogle et al. 2022). These genes involve in transcription and chromatin regulation, RNA processing, splicing and RBP regulation, proteasome and ubiquitin–proteostasis, cell cycle, DNA replication and genome stability, translation initiation and elongation factors, ribosomal proteins, mitochondrial translation, respiratory chain and metabolism, cytoskeleton, motor and structural components, and vesicle trafficking and membrane transport etc. Because each of the essential genes is perturbed directly by CRISPRi in K562 cells, the resulting perturb-seq data provide an interventional framework for estimating causal networks among these genes, representing a context-specific causal effect in this cellular system. While such effects may reflect both direct and mediated regulatory relationships, the proposed approach is expected to yield a biologically interpretable causal gene network among core essential regulators in K562.
@@ 6.1 Evaluation of ARGEN intra-chromosomal causal networks from 3D genome and epigenomic datasets
We first start with analyzing the Perturb-seq data in a chromosome-specific manner and provide evaluations on the causal network learned from the data. We focused on 874 perturbations that had enough number of cells and CRISPRi signal on the targeted genes. For further details on the selection of such genes, refer to Section LABEL:supp-sec:processing_implementation. For each of the 23 human chromosomes, we collected the genes that reside within the same chromosome and applied ARGEN to infer a causal gene network. This resulted in a set of DAGs with around 40 genes and 100 edges per each chromosome on average.
We first implemented an initial sanity check for the directed edges in the DAGs. We checked whether the directional relationships between pairs of genes are indeed causal, instead of being associational. Specifically, as the data is from CRISPRi experiment, where the guide RNA is designed to reduce the expression level of the targeted gene, we would expect to have $\tau_{k}<0$ in the SEM (2). Thus, if the estimated edge coefficient $\hat{\theta}_{jk}<0$ for a parent gene $k$ and a child gene $j$, we would expect the expression level of gene $j$ to increase under an intervention on gene $k$; conversely, if $\hat{\theta}_{jk}>0$, we would expect the expression level of gene $j$ to decrease. We checked whether such a relationship holds for each target gene. As an illustrative example, we collected the entire parents of the gene MCM3 in chromosome 6 (Fig. 3a) and checked the relationship. It can be observed from Fig. 3b that the perturbation on each of the parent genes of MCM3 result in significant increase of MCM3 expression whenever the coefficient $\hat{\theta}_{jk}<0$ and significant decrease of MCM3 expression whenever the $\hat{\theta}_{jk}>0$, aligning with our expectation for CRIPSRi experiments. On the other hand, when MCM3 was perturbed reversely, the expression changes of its parent genes neither exhibit this inverse relationship nor reach comparable levels of significance.
@@
We then checked whether such a result holds for all the edges in the 23 chromosome specific DAGs, and observed that 98% of the ARGEN coefficients $\hat{\theta}_{jk}$ showed negative association with the intervention treatment effect, learned from perturbing a parent of a gene and evaluating the expression change of the gene (Fig. 3e). In comparison, only around 55% of the ARGEN coefficients showed such relationship with the reversed treatment effect, where a child gene was intervened and the change in expression of a parent node of the child was measured, indicating the sign match of nearly random chance (50%). These results reveal that the directionality of the edges is likely to be causal instead of associational.
@@
Next, we checked whether ARGEN DAGs show biological support from 3D genomics and epigenomics. We first investigated the genomic distance pattern of the DAG edges. We specifically gathered the entire $\log_{10}(\mbox{p-values})$ of $\hat{\theta}_{jk}$ and investigated how the edge signal associates with the genomic distance between parent and child genes. We observe that the regulatory elements far apart from the target genes are generally less supported by ARGEN p-values (Fig. 4a), aligning with previous evidences on the genomic distance of gene regulatory elements from the target genes (Fulco et al. 2019).
@@
On top of genomic distance-based evaluation, we further sought for supports from physical proximity information captured from the same system. We specifically explored the actual physical distance structure of the ARGEN edges based on an external Hi-C data of K562 cells (Rao et al. 2014). We first observed that the gene pairs with no Hi-C contact had significantly ($p<2\times 10^{-5}$) lower edge signals than the ones with non-zero Hi-C contacts. In addition, the gene pairs with stronger regulatory signals by ARGEN tend to be supported with higher normalized Hi-C contacts than the ones with less signals ($p<3\times 10^{-6}$, Fig. LABEL:supp-fig:hic_contact).
@@
A/B compartments, which form large genomic territories with relatively high (A) or low (B) gene expression compared to other territories, constitute a class of chromatin domain compartmentalization (Lieberman-Aiden et al. 2009). Consistent with recent findings (Zufferey et al. 2021) that co-regulation within chromatin domains is more prevalent among lowly expressed or less efficiently transcribed genes, as well as evidence that closed chromatin regions exhibit elevated co-methylation across loci (Fortin and Hansen 2015), we observed that the overall edge signal from ARGEN, with compartments learned from Hi-C data, was significantly higher in the B compartment than in the A compartment ($p<5\times 10^{-3}$). Furthermore, among the gene pairs with Transcription Start Site (TSS) distance less than 2 Mega base pairs (Mb), we checked how many of the parent genes are within the same Topologically Associating Domain (TAD), a contiguous regions of genome within which chromatin interactions are enriched (Dixon et al. 2012). We investigated the proportion of such regulatory genes in the gene pairs called as significant by ARGEN’s online FDR procedure in comparison to those in the gene pairs called as not significant. This comparison yielded that those genes within ARGEN significant edges show significantly higher proportion ($p<10^{-3}$) than the other set of genes (Fig. 4c). Furthermore, those regulatory genes called by ARGEN showed significantly higher proportion ($p<10^{-3}$) than the ones called by INSPRE and Naive GLM, while the proportion of INSPRE showed no significant difference compared to that of the gene pairs randomly selected (Fig. 4c).
@@
Finally, we checked whether the gene pairs called from ARGEN showed enriched ChIP-seq signal of binding of transcription factors relevant to K562 and one of the histone marks (H3K27ac) studied in Gasperini et al. (2019), Barry et al. (2024). We specifically quantified the proportion of edges involving enriched ChIP-seq signal in both of the gene promoter regions, defined as 1Kb upstream region from the TSS, and compared the proportion against that of randomly selected gene pairs. We observed that six (yellow colored panels in Fig. 4d) out of eight ChIP-seq targets showed enrichment in the ARGEN edges. In comparison, half of the ChIP-seq targets showed lack of enrichment in the INSPRE edges.